Suppose the market for organically grown wheat is modeled through the following market supply and demand functions:

P = 10 + 0.5QS and P = 22 ‒ 2.5QD,

where QS and QD are in bushels, and P is price per bushel.

What is the consumer surplus? What is the producer surplus?
Answer #1

CS = area of the triangle = 0.5*4*(22-12) = 20

PS = area of the triangle = 0.5*4*(12-10) = 4